OpenStudy (vishweshshrimali5):

If sin x sin y =1/4 and 3tanx=4tany then prove that sin(x+y)=7/16

OpenStudy (vishweshshrimali5):

@hartnn

OpenStudy (vishweshshrimali5):

@satellite73 , @UnkleRhaukus

OpenStudy (vishweshshrimali5):

I was able to get sin(x+y) = 7/12 (cot x) And \[\huge cos x = \cfrac{-1\pm \sqrt{1+36cos^{2}y}}{6 cos y}\]

OpenStudy (vishweshshrimali5):

I have an exam so plsssssssss help me